Consider a capitalist economy which has two sectors: one producing food and the other machine-tools. Assume that capitalists and workers spend their entire earnings on the food. A worker earns $40, 000 in year in either sector. In the food sector, 200 workers work with machines/tools worth $4,000,000 and produce food worth $ 14,000,000 in a year. In the machine/tool sector 100 workers work with machine tools worth $8,000,000 to produce machine/tools worth $18,000, 000 in a year.

  1. What is aggregate demand for food in the economy as a whole:
  2. What is the aggregate demand for machine/ tools in the economy as a whole:
  3. Does the economy meet the condition required for simple reproduction? Explain.
  4. If we now assume that capitalists consume only 1/4 of their profit earnings and reinvest the unconsumed portion (use it to buy new machine tools) what is the aggregate demand for food in the economy as a whole?
  5. What is the aggregate demand for machine tools now (with the assumption of part d)?
Answer #1

Aggregate demand for food as a whole = $40,000*200 + $40,000*100 = $40,000*300 = $12,000,000

Aggregate demand for machine/tools as a whole = $4,000,000 + $8,000,000 = $12,000,000

As the capitalists and workers spend their entire earnings on food, there is no reproduction.

If capitalists consume 1/4th of their earnings, and use the unconsumed portion to buy machine/tools, then aggregate demand for food = 1/4*$12,000,000 = $3,000,000

As 3/4th portion of their earning is used as capital , aggregate demand for machine/tools = $12,000,000+3/4*$12,000,000 = $21,000,000